An executive invests $28,000, some at 7% and some at 4% annual interest. If he receives an annual return of $1720, how much is invested at each rate?
at 7% interest
at 4% interest
please help me solve this. I would be greatly appreciated

Answer by mananth(16949)


Investment I 7.00% per annum $x
Investment II 4.00% per annum $y
x+y=28000 ------------------------1
7.00%x+4.00%y=$1,720.00
Multiply by 100
7x+4y=$172,000.00 --------2
Multiply (1) by -7
we get
-7x-7y= -196000.00
Add this to (2)
-3y=-$24,000.00
divide by -3
y=$8,000.00 investment at 4.00%
Balance $20,000.00 investment at 7.00%
CHECK
20000 @ 7%=$1,400.00
8000 @ 4% = $320.00
TOTAL Interest $1,720.00